OpenStudy (anonymous):

Simplify: 7th root of 3 - the 4th root of 6 + square root of 48 - the square root of 54

OpenStudy (midhun.madhu1987):

\[\sqrt[7]{3} - \sqrt[4]{6} + \sqrt{48} - \sqrt{54}\] Is this the question?

OpenStudy (welshfella):

first factor some of these radical sqrt 48 = sqrt 16 * sqrt 3 = 4 sqrt3 sqrt54 = sqrt 9 * sqrt 6 = 3 sqrt 6